# Secondary spins of extreme mass ratio inspirals: Measurement precision and astrophysical applications

### v1

@@ II.2 Fisher forecast
where the inner product of two signals, $\mathbf{a}(t)$ and $\mathbf{b}(t)$, is defined as Cutler & Flanagan (1994)
@@
and the upper index “*” labels the complex conjugate.
The signal-to-noise ratio (SNR) $\rho$ for a true waveform $\mathbf{h}$ is defined as
@@
For large SNR, the likelihood $\mathcal{L}$ is approximately a Gaussian distribution centered on the true physical values Vallisneri (2008),
@@
is the Fisher information matrix. So, we define the measurement precision through the covariance matrix $\Gamma^{-1}$ by
@@
where dual Riemann tensor $R^{\alpha}_{*\beta\gamma\delta}\equiv\frac{1}{2}R^{\alpha}_{\ \beta\mu\nu}\epsilon^{\mu\nu}_{\ \ \ \gamma\delta}$ and we have included the non-zero components only. When $p^{r}=0$, which corresponds to circular orbits, $Dp^{\phi}/d\tau$ should be 0 as a natural result of its constant azimuthal frequency Han (2010). That is to say, for circular orbits, the momentum evolution and the secondary spin evolution are decoupled, so the waveform should contain little information about $s$ in this situation. This conclusion is consistent with the previous detailed studies Huerta & Gair (2011); Piovano et al. (2021). However, in the case of an eccentric orbit, $p^{r}$ varies and deviates from zero, then the coupling terms appearing in Eq. (42) become nonzero. The effect of the secondary spin is larger for more eccentric orbits as the coupling is proportional to $p^{r}$. Furthermore, for generic orbits, more coupling terms will be released, as $p^{\theta}$ is non-zero and $S^{\mu}$ has more non-zero components, so it is possible to measure the secondary spin with a reasonable accuracy.
@@ III.1 Modeling secondary spins of dry and Hills EMRIs
As briefly summarized in the Introduction, three main EMRI formation channels have been well established in the literature: dry channel Hopman & Alexander (2005); Preto & Amaro-Seoane (2010); Bar-Or & Alexander (2016); Babak et al. (2017b); Amaro-Seoane (2018); Broggi et al. (2022), Hills channel Miller et al. (2005); Raveh & Perets (2021) and wet channel Sigl et al. (2007); Levin (2007); Pan & Yang (2021a); Pan et al. (2021); Pan & Yang (2021b); Pan et al. (2022); Derdzinski & Mayer (2023). A unique signature of wet EMRIs is the low orbital eccentricity in the mHz band. Dry and Hills EMRIs are quite similar in the orbital parameters, therefore how to distinguish them is a remain-to-be-answered question. In the remainder of this section, we will first parametrize the distribution of secondary spins of dry EMRIs and Hills EMRIs, then test whether we can measure the contribution to future EMRI detections of each channel via an injection and inference experiment.
Dry EMRIs keep the birth spins of sBHs formed from individual massive stars, which are of low values as inferred from LVK binary BH merger events Abbott et al. (2019); Roulet & Zaldarriaga (2019); Fuller & Ma (2019); Abbott et al. (2023) and we parameterize the secondary spin distribution in a (semi-)Gaussian form
@@
For comparison, Hills EMRIs keep the birth spins of sBHs formed from binary massive stars. The spins of the first-born sBHs approximately follow the same distribution $P_{1}(s)$ of isolated sBHs, while the spins of the second-born sBHs are largely determined by the tidal spin up during the pre-collapse Helium core phase, thereby follow a different distribution $P_{2}(s)$, which is commonly believed to be a roughly bimodal distribution Zaldarriaga et al. (2018); Bavera et al. (2020); Mandel & Fragos (2020). In the evolution track of a massive star binary, the more massive star evolves faster and collapses into a BH first, and the lighter one later expands largely during its red giant phase, forming an common envelope. Within a dynamical timescale, the binary separation shrinks by orders of magnitude and the common envelope is ejected (see e.g., Ref.  Olejak et al., 2020, for explicit binary evolution examples). In this naked Helium core + first-born BH stage, the He-core can be tidally spun up by the companion BH. In the case of wide binaries, the tidal effect is weak, and the lifetime time of the He-core $T_{\rm He}$ is shorter than the tidal synchronization timescale $T_{\rm sync}$. As a result, the He-core has not been tidally locked at collapse and the spin of the second-born BH scales as Zahn (1977)
@@
where $m_{1}$ and $m_{2}$ are the masses of the first born and second born BHs, respectively. In combination with the minimal binary separation, the He-core radius Kippenhahn & Weigert (1990)
@@ III.2 Population inference basics
The hierarchical Bayesian method has been widely used in population inference of stellar mass binary mergers detected by ground-based detectors LIGO/Virgo/KAGRA. We will use the same method to constrain the EMRI population properties in this subsection. From loud events that can be individually detected $\{\mathbf{d}_{i}\}\ (i=1,...,N_{\rm det})$, one can infer the total number of all events $N_{\rm tot}$ and the population parameters $\mathbf{\Lambda}$, with the population likelihood LIGO Scientific Collaboration & Virgo Collaboration (2019); The LIGO Scientific Collaboration et al. (2021)
@@ IV Conclusion and Discussion
In this paper, we constructed a kludge EMRI waveform model that includes the effect of the secondary spin by numerically integrating the MPD equations (1-3). With this waveform template, we forecast the LISA measurement precision of secondary spins of EMRIs using the Fisher information matrix. As a result, we found that the secondary spin in general is poorly constrained for low-eccentricity and nearly-equatorial EMRIs , but is much better constrained for generic eccentric and/or inclined EMRIs (see Fig. 2), due to stronger coupling between the secondary spin and EMRI orbital dynamics in this case.
@@
With future EMRI detections, the stellar mass BH formation history and the EMRI formation process encoded by the secondary spin can be extracted and used as a probe to stellar evolution and stellar dynamics in nuclear clusters. As a simple example of such astrophysical applications, we did a test on whether dry EMRIs and Hills EMRIs can be distinguished via the secondary spin. Previous studies show that dry and Hills EMRIs are similar in their orbital parameters on the population level, high eccentricities and isotropic inclinations. In this work, we pointed out that the secondary spins of dry and Hill EMRIs are expected to be quite different: the sBH from the dry channel keeps the natal low spin as born in the collapse of a massive star Abbott et al. (2019); Roulet & Zaldarriaga (2019); Fuller & Ma (2019); Abbott et al. (2023), while the sBH from the Hills channel might be spun up during the He-core stage by its companion Zaldarriaga et al. (2018); Bavera et al. (2020); Mandel & Fragos (2020). Therefore, high secondary spins will be a clear signature of Hills EMRIs. As a concrete example, we parameterized the secondary spin distributions of dry and Hills EMRIs as in Eq. (51) and Eq. (52), respectively. Via an injection and inference experiment, we found that several population parameters, including the secondary spin uncertainty $\sigma_{1}$ and the branch ratio $f_{\rm dry}$ of the dry channel, are expected to be well measured from a reasonable number of EMRI detections ($N_{\rm det}=121$). Of course, this conclusion depends on the validity of Eqs. (51,52). On the other hand, if these simple parameterizations are not good approximations to the true distributions, the robust detection of the high-spin peak will still be a clear signature of Hills EMRIs, though we may not trust the exact interpretation of individual population parameters in this case.
@@
Wet EMRIs are expected to be of low eccentricity in the LISA sensitivity band due to the efficient eccentricity damping by density waves as migrating along the accretion disk. If AGN accretion is coherent, wet EMRIs that are formed in the AGN disk are preferentially on the equator. In this case, the secondary spins are poorly constrained (see Fig. 2). If AGN accretion is chaotic with gas feeding from a random direction in each accretion episode, wet EMRIs that are brought in by the outer disk are expected to be misaligned with the inner equatorial disk Lyu et al. (2024). In this case, the secondary spins could be constrained (Fig. 2) and used as probe to the accretion history of EMRIs embedded in the AGN disk Pan & Yang (2021b); Li et al. (2022); Chen & Lin (2023).

### v3

@@ II.2 Fisher forecast
where the inner product of two signals, $\mathbf{a}(t)$ and $\mathbf{b}(t)$, is defined as Cutler and Flanagan (1994)
@@
and the upper index “*” labels the complex conjugate. And the signal-to-noise ratio (SNR) $\rho$ for a true waveform $\mathbf{h}$ is defined as
@@
To estimate the system parameter, one should fully explore the posterior distribution by Bayesian analysis (e.g., Markov-chain Monte Carlo). But for EMRI waveforms, to search across the large parameter space requires template waveforms that can be generated in less than a second, which remains unachieved for the generic orbit and lies beyond the study’s primary objectives. So, in this paper, we use the Fisher Information Matrix (FIM) approximation as a practical method to evaluate the measurement precision, which is widely adopted in GW research, although with known inadequacies Rodriguez et al. (2013); Vallisneri (2008). For large SNR, the likelihood $\mathcal{L}$ is approximately a Gaussian distribution centered on the true physical values Vallisneri (2008),
@@
is the FIM. So, we define the measurement precision through the covariance matrix $\Gamma^{-1}$ by
@@
where dual Riemann tensor $R^{\alpha}_{*\beta\gamma\delta}\equiv\frac{1}{2}R^{\alpha}_{\ \beta\mu\nu}\epsilon^{\mu\nu}_{\ \ \ \gamma\delta}$ and we have included the non-zero components only. When $p^{r}=0$, which corresponds to circular orbits, $Dp^{\phi}/d\tau$ should be 0 as a natural result of its constant azimuthal frequency Han (2010). That is to say, for circular orbits, the momentum evolution and the secondary spin evolution are decoupled, so the waveform should contain little information about $s$ in this situation. This conclusion is consistent with the previous detailed studies Huerta and Gair (2011); Piovano et al. (2021). However, in the case of an eccentric orbit, $p^{r}$ varies and deviates from zero, then the coupling terms appearing in Eq. (44) become nonzero. The effect of the secondary spin is larger for more eccentric orbits as the coupling is proportional to $p^{r}$. Furthermore, for generic orbits, more coupling terms will be released, as $p^{\theta}$ is non-zero and $S^{\mu}$ has more non-zero components, so it is possible to measure the secondary spin with a reasonable accuracy.
@@ III.1 Modeling secondary spins of dry and Hills EMRIs
As briefly summarized in the Introduction, three main EMRI formation channels have been well established in the literature: dry channel Hopman and Alexander (2005); Preto and Amaro-Seoane (2010); Bar-Or and Alexander (2016); Babak et al. (2017b); Amaro-Seoane (2018); Broggi et al. (2022), Hills channel Miller et al. (2005); Raveh and Perets (2021) and wet channel Sigl et al. (2007); Levin (2007); Pan and Yang (2021a); Pan et al. (2021); Pan and Yang (2021b); Pan et al. (2022); Derdzinski and Mayer (2023). A unique signature of wet EMRIs is the low orbital eccentricity in the mHz band. Dry and Hills EMRIs are quite similar in the orbital parameters, therefore how to distinguish them is a remain-to-be-answered question. In the remainder of this section, we will first parametrize the distribution of secondary spins of dry EMRIs and Hills EMRIs, then test whether we can measure the contribution to future EMRI detections of each channel via an injection and inference experiment.
Dry EMRIs keep the birth spins of sBHs formed from individual massive stars, which are of low values as inferred from LVK binary BH merger events Abbott et al. (2019); Roulet and Zaldarriaga (2019); Fuller and Ma (2019); Abbott et al. (2023) and we parameterize the secondary spin distribution in a (semi-)Gaussian form
@@
For comparison, Hills EMRIs keep the birth spins of sBHs formed from binary massive stars. The spins of the first-born sBHs approximately follow the same distribution $P_{1}(s)$ of isolated sBHs, while the spins of the second-born sBHs are largely determined by the tidal spin up during the pre-collapse Helium core phase, thereby follow a different distribution $P_{2}(s)$, which is commonly believed to be a roughly bimodal distribution Zaldarriaga et al. (2018); Bavera et al. (2020); Mandel and Fragos (2020); Adamcewicz et al. (2024). In the evolution track of a massive star binary, the more massive star evolves faster and collapses into a BH first, and the lighter one later expands largely during its red giant phase, forming an common envelope. Within a dynamical timescale, the binary separation shrinks by orders of magnitude and the common envelope is ejected (see e.g., Ref.  Olejak et al., 2020, for explicit binary evolution examples). In this naked Helium core + first-born BH stage, the He-core can be tidally spun up by the companion BH. In the case of wide binaries, the tidal effect is weak, and the lifetime time of the He-core $T_{\rm He}$ is shorter than the tidal synchronization timescale $T_{\rm sync}$. As a result, the He-core has not been tidally locked at collapse and the spin of the second-born BH scales as Zahn (1977)
@@
where $m_{1}$ and $m_{2}$ are the masses of the first born and second born BHs, respectively. In combination with the minimal binary separation, the He-core radius Kippenhahn and Weigert (2012)
@@ III.2 Population inference basics
The hierarchical Bayesian method has been widely used in population inference of stellar mass binary mergers detected by ground-based detectors LIGO/Virgo/KAGRA. We will use the same method to constrain the EMRI population properties in this subsection. From loud events that can be individually detected $\{\mathbf{d}_{i}\}\ (i=1,...,N_{\rm det})$, one can infer the total number of all events $N_{\rm tot}$ and the population parameters $\mathbf{\Lambda}$, with the population likelihood LIGO Scientific Collaboration and Virgo Collaboration (2019); The LIGO Scientific Collaboration et al. (2021)
@@ IV Conclusion and Discussion
In this paper, we constructed a kludge EMRI waveform model that includes the effect of the secondary spin by numerically integrating the MPD equations (3-5) and PN radiation reactions (15a). With this waveform template, we forecast the LISA measurement precision of secondary spins of EMRIs using the Fisher information matrix. As a result, we found that the secondary spin in general is poorly constrained for low-eccentricity and nearly-equatorial EMRIs, but is much better constrained for generic eccentric and/or inclined EMRIs (see Fig. 2), due to stronger coupling between the secondary spin and EMRI orbital dynamics in this case. Note that we certainly won’t expect this rudimentary template to achieve the requisite precision for future waveform-matched filtering, and the insufficiently verified use of Fisher-Matrix analysis in this paper may also introduce uncertainties to the quantitative results.
@@
With future EMRI detections, the stellar mass BH formation history and the EMRI formation process encoded by the secondary spin can be extracted and used as a probe to stellar evolution and stellar dynamics in nuclear clusters. As a simple example of such astrophysical applications, we did a test on whether dry EMRIs and Hills EMRIs can be distinguished via the secondary spin. Previous studies show that dry and Hills EMRIs are similar in their orbital parameters on the population level, high eccentricities and isotropic inclinations. In this work, we pointed out that the secondary spins of dry and Hill EMRIs are expected to be quite different: the sBH from the dry channel keeps the natal low spin as born in the collapse of a massive star Abbott et al. (2019); Roulet and Zaldarriaga (2019); Fuller and Ma (2019); Abbott et al. (2023), while the sBH from the Hills channel might be spun up during the He-core stage by its companion Zaldarriaga et al. (2018); Bavera et al. (2020); Mandel and Fragos (2020). Therefore, high secondary spins will be a clear signature of Hills EMRIs. As a concrete example, we parameterized the secondary spin distributions of dry and Hills EMRIs as in Eq. (53) and Eq. (54), respectively. Via an injection and inference experiment, we found that several population parameters, including the secondary spin uncertainty $\sigma_{1}$ and the branch ratio $f_{\rm dry}$ of the dry channel, are expected to be well measured from a reasonable number of EMRI detections ($N_{\rm det}=121$). Of course, this conclusion depends on the validity of Eqs. (53,54). On the other hand, if these simple parameterizations are not good approximations to the true distributions, the robust detection of the high-spin peak will still be a clear signature of Hills EMRIs, though we may not trust the exact interpretation of individual population parameters in this case.
@@
Wet EMRIs are expected to be of low eccentricity in the LISA sensitivity band due to the efficient eccentricity damping by density waves as migrating along the accretion disk. If AGN accretion is coherent, wet EMRIs that are formed in the AGN disk are preferentially on the equator. In this case, the secondary spins are poorly constrained (see Fig. 2). If AGN accretion is chaotic with gas feeding from a random direction in each accretion episode, wet EMRIs that are brought in by the outer disk are expected to be misaligned with the inner equatorial disk Lyu et al. (2024). In this case, the secondary spins could be constrained (Fig. 2) and used as probe to the accretion history of EMRIs embedded in the AGN disk Pan and Yang (2021b); Li et al. (2022); Chen and Lin (2023).